# Optimal Execution with Reinforcement Learning

### v1

@@ V Reinforcement learning
A discrete-time Markov Decision Process (MDP) [Puterman, 1990] is defined as a tuple $\langle\mathcal{S},\mathcal{A},\mathbb{P},\mathcal{R},\gamma,\mu\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathbb{P}(\cdot|s,a)$ is a Markovian transition model that assigns to each state-action pair $(s,a)$ the probability of reaching the next state $s^{\prime}$, $\mathcal{R}(s,a)$ is a bounded reward function, $\gamma\in[0,1)$ is the discount factor, and $\mu$ is the distribution of the initial state. The policy of an agent is characterized by $\pi(\cdot|s)$, which defines for each state $s$ an action with a probability distribution over the action space.
@@
The objective in RL is the maximization of the expected return, given an initial state distribution:
$J_{\pi}\coloneqq\underset{s_{0}\sim\mu}{\mathbb{E}_{\pi}}\left[\sum_{i=1}^{\tau}\gamma^{i-1}r_{i}.\right].$
Where the return is the discounted sum of the rewards.
@@
1:Initialize replay memory $D$ to capacity $N$
@@
Numerous RL algorithms are available [Schulman et al., 2017b, Schulman et al., 2015, Mnih et al., 2015]. In this work, we focus on the Deep Q-Network (DQN) algorithm, a model-free, online, off-policy reinforcement learning approach, as it provided the best learning results. The DQN algorithm is detailed in algorithm 1.
V-A Embedding Optimal Execution as an MDP
  • State: Percentage holdings remaining, percentage time remaining, volume imbalance up to 5 levels of the limit order book, best bid price, best ask price.
  • Actions: 5 possibile actions: do nothing, consume $Q_{t}^{k}=Q_{min}\times k$, $k=1,\dots,4$.
  • Reward:
    $r_{t}=\underbrace{Q_{t}^{k}\times(P_{0}-P_{t})}_{\text{implementation shortfall}}-\underbrace{\alpha d_{t}}_{\text{penalty}},$ (5)
    where $P_{t}$ is the average execution price, $P_{0}$ the arrival price, $d_{t}$ is the depth consumed by the market order at time $t$, and $\alpha>0$. Once the execution ends, even if the available time is still not over, the reward will be 0, while if the execution has not ended by the end of the execution period, there will be an additional large penalty.
V-B Tailoring the environment
Although we have the standard objective in Equation 2, how to appropriately define the MDP to achieve that objective is not straightforward. In fact, there are several factors which make that objective challenging:
  1. 1.
  2. 2.
  3. 3.
To tackle the first factor, we included the penalty in the reward function, which has the objective of emphasizing that we want to avoid market impact. To address the second, we include an additional large penalty if the execution is not finished by the end of the total available period. To address the third factor, the penalty in the reward function is enabling. In fact, thanks to the penalty, the reward is negative in most cases. As a reward of 0 obtained consistently once the execution is finished is preferable to a negative one reward, the learning agent should be incentivized to finish earlier.
We also explored different state and action space setups, and those described in Section V-A gave the best results.
VI Experimental Results
@@
RL algorithm setup
We ran a hyperparameter tuning which gave the following optimal parameters. We employed DQN, with a neural network architecture that features fully connected $2$ input layers composed of $50$ and $20$ neurons. The learning rate schedule implements a linear decrease from $10^{-3}$ to 00 in $90,000$ steps, while the exploration rate follows a $\epsilon$Greedy search from $1$ to $0.02$ in $10,000$ steps. In addition, we implement a state history length of $4$ and a market data buffer of length $50$. This approach is aimed at ensuring the stability of DQN. Finally, the objective function is a discounted sum of rewards with a discount factor of $\gamma=0.9999$.
Baseline algorithms
@@
  • TWAP algorithm: defined in equation 4.
  • Passive algorithm: this policy mostly keeps the agent inactive, with a 60% chance of doing nothing. Occasionally, with a 40% chance, it randomly chooses and executes a quantity among four available options, each with equal likelihood.
  • Aggressive algorithm: purchase a fixed amount $2\times Q_{\text{min}}$ each second.
  • Random algorithm: sample whether to use a market order or do nothing (quantity of 0). There is a 50% chance of doing nothing. Alternatively, it randomly selects and executes one of three actions with equal probability: do nothing, or consume $Q_{t}^{k}=Q_{min}\times k$, with $k=1,\dots,3$.
VI-A Experiments
@@
In Figure 4, we evaluate out-of-sample the optimized reinforcement learning policy and compare it to the three baselines. The distribution is created after running 20, $30$-minute windows with different seeds. We observe the average step implementation shortfall (see Equation 5). In this case we see that the RL agent generates a lower variance compared to the baselines with also a higher average. This can be interpreted as the fact that the RL execution strategy is capable of executing consistently close to the arrival price while minimizing market impact.
Figure 5: Ask prices during the RL execution.
Figure 5 presents the ask prices during execution by the RL agent over time. The stability of ask prices under the RL policy indicates the agent’s ability to minimize market impact and adverse price movements effectively.
Figure 6: Execution trajectory of the RL agent: amount executed over time.
Figure 6 shows the execution trajectory of the RL agent: how much it has executed of the total size. The agent executes a significant portion of its holdings rapidly, followed by a more controlled and steady approach as the execution period progresses. This pattern aligns with optimal execution principles, balancing the trade-off between immediate market impact and long-term price stability. By front-loading the execution and then tapering off, the agent minimizes the risk of adverse market movements while efficiently managing its inventory.
Figure 7: Bid and ask prices along with actions taken by the RL agent over time.
Figure 7 provides a detailed view of the bid and ask prices alongside the actions executed by the RL agent at various time steps. The figure illustrates that the actions are aligned with a minimal deviation between the bid and ask prices, indicating effective control over the execution strategy without causing market disruption.
Figure 8: Distribution of the spreads generated by the different strategies.
In Figure 8, we observe the distribution of price spreads generated by the different trading strategies. We can see that the RL widens a bit the spreads compared to the TWAP, but this is because it finishes the execution earlier as we saw in Figure 6. If we consider instead the aggressive strategy, it generates very wide spreads and so was taken out of the plot in order to visualize better the other policies.
Figure 9: Distribution of the volume imbalance generated by the different strategies.
Figure 9 shows the distribution of the volume imbalance generated by the RL agent’s execution. A market impact would be visible with a skewed imabalance. In this case the imbalance is centered, indicating a consistent strategy for maintaining balance in the LOB.
Metric RL TWAP Passive Random
Imp. Short. -0.022 -0.028 -0.024 -0.049
Avg. Pen. -0.001 $0$0.0035 -0.003
Time % 0.451 0.998 0.727 0.553
TABLE I: Comparison of normalized performance metrics across different trading strategies.
@@
VII Conclusions
The results of this study demonstrate the effectiveness of using reinforcement learning (RL) for optimal execution in financial markets. Using the deep Q-Network (DQN) algorithm within a simulated market environment, we observed significant improvements in execution performance compared to the chosen benchmark strategies.
Our experiments show that the RL agent consistently achieved higher returns and lower variance in implementation shortfall. The agent’s ability to adapt to market conditions and execute trades close to the arrival price resulted in minimized market impact and transaction costs. This stability is crucial to maintain favorable execution conditions.
The execution trajectory analysis highlighted the strategic balance of the RL agent between immediate market impact and long-term price stability. The agent effectively managed its holdings, executing a significant portion rapidly at the beginning and then proceeding more steadily. This approach is aligned with the principles of optimal execution, ensuring efficient inventory management.
Although our findings are promising, it is important to acknowledge areas for future exploration to enhance the robustness and applicability of RL-based strategies. The simulated market environment, though realistic, can be further refined to capture a wider array of market dynamics and participant behaviors. Expanding the range of market conditions under which the RL models are tested will provide deeper insights into their performance and adaptability.
Furthermore, the computational resources required for training RL models are considerable, and optimizing these processes will be crucial for practical implementation. Streamlining the training process and improving the efficiency of the algorithms will facilitate their broader application in real-world scenarios.
In conclusion, the positive results of our study underscore the potential of reinforcement learning to address the complexities of financial markets. This approach provides traders and institutions with enhanced tools for efficient and effective execution. Continued research and development will be essential to ensure the robustness and adaptability of these models, paving the way for their successful application in dynamic and competitive trading environments.

### v2

@@ IV Reinforcement Learning
A discrete-time Markov Decision Process (MDP) (see [27]) is defined as a tuple $\langle\mathcal{S},\mathcal{A},\mathbb{P},\mathcal{R},\gamma,\mu\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $\mathbb{P}(\cdot|s,a)$ is a Markovian transition model that assigns to each state-action pair $(s,a)$ the probability of reaching the next state $s^{\prime}$, $\mathcal{R}(s,a)$ is a bounded reward function, $\gamma\in[0,1[$ is the discount factor, and $\mu$ is the distribution of the initial state. The policy of an agent is characterized by $\pi(\cdot|s)$, which defines for each state $s$ an action with a probability distribution over the action space.
IV-A Embedding Optimal Execution as an MDP
  • State space: Percentage holdings remaining, percentage time remaining, volume imbalance up to 5 levels of the LOB, best bid price, best ask price.
  • Action space: 5 possibile actions: do nothing, consume $Q_{t}^{k}=Q_{min}\times k$, $k=1,\dots,4$.
  • Reward function:
    $r_{t}=\underbrace{Q_{t}^{k}\times(P_{0}-P_{t})}_{\text{implementation shortfall}}-\underbrace{\alpha d_{t}}_{\text{penalty}}-\underbrace{\beta I_{T}\mbox{1\hskip-2.5ptl}_{\{t=T\}}}_{\text{terminal penalty}},$ (4)
    where $P_{t}$ is the average execution price, $P_{0}$ the arrival price, $d_{t}$ is the depth consumed by the market order at time $t$, and $\alpha>0$, $I_{T}$ is the remaining inventory at the end of the execution period and $\beta>0$. Once the execution is completed, the reward remains zero for the rest of the period, whereas any residual inventory at time $T$ incurs a terminal penalty.
IV-B Tailoring the Environment
Although we have the standard objective in Equation (1), how to appropriately define the MDP to achieve that objective is not straightforward. In fact, there are several factors which make that objective challenging:
  1. 1.
  2. 2.
  3. 3.
To tackle the first factor, we included the penalty in the reward function, which has the objective of emphasizing that we want to avoid market impact. To address the second, we include an additional large penalty if the execution is not finished by the end of the total available period. To address the third factor, the penalty in the reward function is enabling. In fact, thanks to the penalty, the reward is negative in most cases. As a 0 reward obtained consistently once the execution is finished is preferable to a negative reward, the learning agent should be incentivized to finish earlier.
We also explored different state and action space setups, and those described in Section IV-A gave the best results.
IV-C Algorithm
@@
The objective in RL is the maximization of the expected return, given an initial state distribution
$J_{\pi}:=\underset{s_{0}\sim\mu}{\mathbb{E}_{\pi}}\Big[\sum_{i=1}^{\tau}\gamma^{i-1}r_{i}.\Big],$
where the return is the discounted sum of the rewards. Numerous RL algorithms are available (see [28, 29, 30]). In this work, we focus on the Deep Q-Network (DQN) algorithm, a model-free, online, off-policy reinforcement learning approach, as it provided the best learning results. The DQN algorithm is detailed in algorithm 1.
@@
1:Initialize replay memory $D$ to capacity NN
@@
V Experimental Results
@@ V-A Environment setup
V-B RL algorithm setup
We ran a hyperparameter tuning which gave the following optimal parameters. We employed DQN, with a neural network architecture that features fully connected $2$ input layers composed of $50$ and $20$ neurons. The learning rate schedule implements a linear decrease from $10^{-3}$ to 0 in $90,000$ steps, while the exploration rate follows a $\epsilon$-Greedy search from $1$ to $0.02$ in $10,000$ steps. In addition, we implement a state history length of $4$ and a market data buffer of length $50$. This approach is aimed at ensuring the stability of DQN. Finally, the objective function is a discounted sum of rewards with a discount factor of $\gamma=0.9999$. This means that the agent treats all future rewards almost as equally important which is useful in execution problems where rewards come at the end (e.g., penalizing slippage at end of episode).
We also experimented with alternative state spaces (continuous and discrete) and algorithms such as TRPO, PPO, and actor–critic, but DQN consistently proved the most reliable.
V-C Baseline algorithms
@@
  • Time-Weighted Average Price (TWAP): defined in equation (3), corresponding to the risk-neutral solution of the Almgren–Chriss model.
  • Passive Policy (PP): this policy mostly keeps the agent inactive, with a 60% chance of doing nothing. With a 40% chance, it randomly chooses and executes a quantity among four available options, each with equal likelihood.
  • Random Policy (RP): sample whether to use a market order or do nothing (quantity of 0). There is a 50% chance of doing nothing. Alternatively, it randomly selects and executes one of three actions with equal probability: do nothing, or consume $Q_{t}^{k}=Q_{min}\times k$, with $k=1,\dots,3$.
V-D Experiments
@@ V-D1 Standard Configuration Analysis
In Figure 5, we evaluate out-of-sample the optimized reinforcement learning policy and compare it to the three baselines. The distribution is created after running 20, $30$-minute windows with different seeds. We observe the average step implementation shortfall (see Equation (4)). In this case we see that the RL agent generates a lower variance compared to the baselines with also a higher average. This suggests that the RL execution strategy is capable of executing consistently close to the arrival price while minimizing market impact.
(a) Actions taken by the RL agent for a single trajectory over time.
(b) Execution trajectory of the RL agent: amount executed over time.
(c) Temporal evolution of Q-values with respect to actions.
Figure 6: RL agent behavior in a simulated execution task
Figure 6b shows the execution trajectory of the RL agent: how much it has executed of the total size. The agent executes a significant portion of its holdings rapidly, followed by a more controlled and steady approach as the execution period progresses. This pattern aligns with optimal execution principles, balancing the trade-off between immediate market impact and long-term price stability. By front-loading the execution and then tapering off, the agent minimizes the risk of adverse market movements while efficiently managing its inventory. Figure 6a provides a detailed view of the bid and ask prices alongside the actions executed by the RL agent at various time steps. The figure illustrates that the actions are aligned with a minimal deviation between the bid and ask prices, indicating effective control over the execution strategy without causing market disruption.
(a) Distribution of the spreads.
(b) Distribution of the volume imbalance.
Figure 7: Empirical distributions of spreads and imbalance for different strategies.
In Figure 7a, we observe that the RL strategy executes with very tight spreads, indicating that it trades at moments when liquidity is highest in the limit order book, contrarily to the other strategies. Figure 7b shows the distribution of the volume imbalance generated by the RL agent’s execution. A market impact would be visible with a skewed imabalance. In this case the imbalance is centered, indicating a consistent strategy for maintaining balance in the LOB.
Strategy $\mathbb{E}$(IS) $\mathbb{E}$(Pen) $\mathbb{E}$(T) $\sigma^{2}$(IS)
RL 0.0002 -0.005 0.63 0.0011
TWAP -0.0321 0.000 0.998 0.0014
Passive -0.0106 -0.007 0.77 0.003
Random -0.016 -0.003 0.576 0.001
Table I: Comparison of normalized performance metrics with $1000$ noise agents and $12$ momentum agents.
@@
V-D2 Robustness to Market Agent Configurations
Next, we demonstrate the robustness of our RL strategy by showing that it consistently outperforms all benchmark approaches in terms of both the expected implementation shortfall and the variance of performance across different market configurations. In particular, Figures 8 and Tables II and III show the performance under varying numbers of noise agents and momentum agents.
(a) 10 noise agents
(b) 1000 noise agents
(c) 2000 noise agents
(d) 6 momentum agents
(e) 12 momentum agents
(f) 24 momentum agents
Figure 8: Implementation shortfall normalized by initial order size. Top row varies momentum agents with $12$ fixed. Bottom row varies noise agents with $1000$ fixed.
Strategy #NA $\mathbb{E}$(IS) $\mathbb{E}$(Pen) $\mathbb{E}$(T) $\sigma^{2}$(IS)
RL 10 -0.017 -0.0001 0.55 0.0039
2000 0.0173 -0.0001 0.56 0.00043
TWAP 10 -0.067 0.000 0.998 0.0022
2000 -0.0184 -0.0035 0.998 0.0024
Random 10 -0.026 -0.003 0.563 0.0004
2000 -0.039 -0.002 0.61 0.00044
Passive 10 -0.048 -0.035 0.7 0.0024
2000 -0.0076 -0.003 0.753 0.003
Table II: Comparison of normalized performance metrics with varying noise agents and $12$ momentum agents.
Strategy #MA $\mathbb{E}$(IS) $\mathbb{E}$(Pen) $\mathbb{E}$(T) $\sigma^{2}$(IS)
RL 6 -0.022 -0.001 0.916 0.0038
24 0.004 -0.001 0.51 0.0035
TWAP 6 -0.028 0.000 0.998 0.0068
24 -0.049 -0.0015 0.998 0.024
Random 6 -0.049 -0.002 0.573 0.0013
24 -0.036 0.000 0.54 0.024
Passive 6 -0.024 -0.001 0.79 0.0031
24 0.0014 -0.0021 0.71 0.049
Table III: Comparison of normalized performance metrics with $1000$ noise agents and varying momentum agents.
We test whether RL outperforms the benchmark strategies on implementation shortfall (IS) using one-sided Student’s two-sample $t$-tests with pooled variance (per-strategy $n=50$). The hypotheses are
$H_{0}:\,\mathbb{E}(\text{IS})_{\text{RL}}\,\leq\,\mathbb{E}(\text{IS})_{S}\quad\text{vs.}\quad H_{1}:\,\mathbb{E}(\text{IS})_{\text{RL}}\,>\,\mathbb{E}(\text{IS})_{S},$
where $S\in\{\text{TWAP},\;\text{Random},\;\text{Passive}\}$ denotes the benchmark strategy. Decisions are made against the one-sided 5% critical value $t_{0.95}(98)\approx 1.660$.
With $10$ noise agents, RL significantly outperforms TWAP ($t=4.52$) and Passive ($t=2.08$). With $2000$ noise agents, RL significantly outperforms all benchmarks: TWAP ($t=4.75$), Random ($t=8.49$), and Passive ($t=2.42$). With $6$ momentum agents, RL significantly outperforms Random ($t=2.14$). With $24$ momentum agents, RL significantly outperforms TWAP ($t=2.26$) and Random ($t=1.81$).
Overall, RL consistently achieves statistically significant gains in the baseline and under noise-agent stress, and also demonstrates advantages in momentum-driven environments.
VI Conclusions
The results of this study demonstrate the effectiveness of using reinforcement learning for optimal execution in financial markets. Using the DQN algorithm within a simulated market environment, we observed significant improvements in execution performance compared to the chosen benchmark strategies. Our experiments show that the RL agent consistently achieved higher returns and lower variance in implementation shortfall. The agent’s ability to adapt to market conditions and execute trades close to the arrival price resulted in minimized market impact and transaction costs. The execution trajectory analysis highlighted the strategic balance of the RL agent between immediate market impact and long-term price stability. The agent managed its inventory efficiently, executing a large portion early before transitioning to a steadier pace. This behavior aligns with optimal execution principles and reflects effective inventory control.
Although our findings are promising, future work could use more tractable simulators such as zero-intelligence frameworks that capture key order book dynamics. This would provide a clearer benchmark to assess RL learning capabilities under a wider range of market conditions.